You work in a materials testing lab and your boss tells you to increase the temperature of a sample by 31.5 ∘C . The only thermometer you can find at your workbench reads in degrees Fahrenheit.

(A)

If the initial temperature of the sample is 68.5 ∘F, what is its temperature in degrees Fahrenheit when the desired temperature increase has been achieved?

Express your answer in degrees Fahrenheit to three significant figures.
Answer #1

Initial temperature in Fahrenheit = Tf1 = 68.5 oF

Initial temperature in Celsius = Tc1

Tf1 = 1.8Tc1 + 32

68.5 = 1.8Tc1 + 32

Tc1 = 20.278 oC

Final temperature in Fahrenheit = Tf2

Final temperature in Celsius = Tc2

Increase in temperature of the sample in Celsius = Tc = 31.5 oC

Tc = Tc2 - Tc1

31.5 = Tc2 - 20.278

Tc2 = 51.778 oC

Tf2 = 1.8Tc2 + 32

Tf2 = (1.8)(51.778) + 32

Tf2 = 125.2 oF

Rounding off to two three significant figures,

Tf2 = 125 oF

A) Temperature in Fahrenheit when the desired temperature increase has been achieved = 125.2 oF